OpenStudy (anonymous):

I'll give a medal to whoever explains it best. The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (zpupster):

4(1) + 1 = ? 4(2) + 1 = ? 4(3) + 1 = ? ...

OpenStudy (anonymous):

5 9 13

OpenStudy (anonymous):

Oh wow, ok I'm stupid

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

can I ask you another @zpupster ?